Christopher Columbus has just spent 20 pieces of silver at the mariner's mall. he bought a hat, a cape, and a spyglass. The spyglasses cost 5 more pieces of silver than the cape. The cape cost 3 more pieces of silver than the hat. how much did each item cost.
Answer by TimothyLamb(4379)

x = pieces of silver for hat
y = pieces of silver for cape
z = pieces of silver for spyglass
---
x + y + z = 20
z = y + 5
y = x + 3
---
put the system of linear equations into standard form
---
x + y + z = 20
y - z = -5
x - y = -3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pieces of silver for hat = 3
y = pieces of silver for cape = 6
z = pieces of silver for spyglass = 11
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---